# Many-Body Rashba Spin-Orbit Interaction and Exciton Spin Relaxation in Atomically Thin Semiconductor Structures

### v1

@@ Abstract
We propose a previously unexplored spin-orbit interaction mechanism by establishing a mesoscopic many-particle Rashba Hamiltonian. In lowest order, this Hamiltonian self-consistently describes exciton spin relaxation in monolayer transition metal dichalcogenides (TMDC) due to local electric fields caused by spatial asymmetries in the dielectric environment. For a monolayer MoSe2 on a SiO2 substrate above 77 K showing a meV bright-dark splitting, the local electric field causes fast intravalley spin relaxation on a sub-picosecond timescale, whereas it is negligible for other TMDCs with larger bright-dark splitting.
@@ I Introduction
Atomically thin semiconductors Wang et al. (2018a) host a broad variety of incoherent exciton scattering mechanisms. Based on the valley-selective optical selection rules, next to spin-conserving exciton-phonon scattering Selig et al. (2018); Brem et al. (2020); Wallauer et al. (2021); Katzer et al. (2023), Coulomb scattering mechanisms such as spin-conserving Dexter interaction Berghäuser et al. (2018) and spin-nonconserving intra- and intervalley exchange Guo et al. (2019); Maialle et al. (1993); Yu and Wu (2014); Selig et al. (2019, 2020); Kwong et al. (2021) via double spin flips are relevant. Moreover, single spin flips can occur via, e.g., Rashba interaction. Rashba spin-orbit interaction can already affect linear optical spectra of a spin-forbidden mixture of $s$-$p$ orbital excitons in monolayers Cao et al. (2024, 2025). However, these signatures are very weak and have not been measured until now. For $s$-orbital excitonic states, which are usually measured in absorption or photoluminescence due to their large optical oscillator strength, no brightening of spin-dark excitonic states in out-of-plane electric fields can occur, since Rashba spin coupling vanishes at zero center-of-mass (COM) momentum within the light cone Mittenzwey et al. (2025). Thus, Rashba spin coupling is only relevant for incoherent $s$-orbital excitons occupying COM momenta well beyond the light cone. In contrast to Coulomb-exchange-induced double spin flips, single spin flip scattering mechanisms are less understood, as they are often included via external parameters Ochoa and Roldán (2013); Dery and Song (2015); Song and Dery (2013); Yang et al. (2020) or resulting from a complex band structure Wang et al. (2018b); Molina-Sánchez et al. (2017). Recently, chiral phonons have been identified as an important single spin-flip mechanism Lagarde et al. (2024); Chan et al. (2025); Zhang and Cao (2022).
In this work, we derive a new many-body Rashba spin-orbit interaction mechanism Winkler (2003); Bychkov and Rashba (1984); Kormányos et al. (2014):
@@
which results from self-consistently described out-of-plane electric fields $\mathbf{E}(\mathbf{r},z)$, induced by the dielectric environment. Here, $\bm{\sigma}=\begin{pmatrix}\sigma_{x}&\sigma_{y}&\sigma_{z}\end{pmatrix}^{\top}$ is the vector of the Pauli matrices $\sigma_{i}$, $\mathbf{p}=\frac{\hbar}{\mathrm{i}}\nabla$ is the momentum operator and $\alpha_{\text{BR}}$ is the Rashba coupling constant, which, in solids, is greatly enhanced compared to the atomic constant appearing in the Pauli equation Kormányos et al. (2014); Winkler (2003).
Starting from Eq. (1), we derive a second-quantized description of the Rashba interaction induced by the optically induced (image) charge density due to spatial asymmetries in the dielectric environment of an atomically thin semiconductor (Sec. II), which is subsequently transformed into an excitonic description (Sec. III). In combination with exciton-phonon interaction generating non-zero COM-momentum excitons (Sec. IV), the described spin-orbit interaction leads to a phonon-assisted exciton spin relaxation. We illustrate the new mechanism for a monolayer MoSe2 and quantify its impact on the incoherent exciton dynamics via microscopic calculations (Sec. V).
@@
$\begin{split}&\hat{H}_{\text{BR}}=\sum_{\begin{subarray}{c}\mathbf{k},\mathbf{k}^{\prime},\mathbf{q},\\ \lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}\\ \xi,\xi^{\prime},\xi^{\prime\prime},s,s^{\prime}\end{subarray}}E_{z,\mathbf{q}+\mathbf{K}^{\xi^{\prime}}-\mathbf{K}^{\xi}}^{\text{loc}}F_{\lambda_{2},\lambda_{3},\mathbf{k}^{\prime}-\mathbf{q},\mathbf{k}^{\prime}}^{\xi^{\prime\prime}+\xi-\xi^{\prime},\xi^{\prime\prime},s^{\prime},s^{\prime}}\\ &\,\quad\times S_{\mathbf{k}+\mathbf{q},\mathbf{k}}^{\lambda_{1},\lambda_{4},\xi^{\prime},\xi,\bar{s},s}{\hat{a}^{\dagger}}{\vphantom{a}}_{\lambda_{1},\mathbf{k}+\mathbf{q}}^{\xi^{\prime},\bar{s}}{\hat{a}^{\dagger}}{\vphantom{a}}_{\lambda_{2},\mathbf{k}^{\prime}-\mathbf{q}}^{\xi^{\prime\prime}+\xi-\xi^{\prime},s^{\prime}}{\hat{a}^{\vphantom{\dagger}}}{\vphantom{a}}_{\lambda_{3},\mathbf{k}^{\prime}}^{\xi^{\prime\prime},s^{\prime}}{\hat{a}^{\vphantom{\dagger}}}{\vphantom{a}}_{\lambda_{4},\mathbf{k}}^{\xi,s}.\end{split}$ (9)
@@
where $\epsilon_{0}$ is the vacuum permittivity and $\epsilon_{s,\bot}$ is the relative out-of-plane dielectric constant of the thin semiconductor. The unitless functions $g_{\mathbf{q}}$ and $f_{\mathbf{q}}$, which encode the momentum dependence determined by the dielectric structure, are given in Eq. (41) and Eq. (42) in App. B, respectively. The electric field in Eq. (10) stems solely from the homogeneous solution of the Poisson equation, which arises due to image charges determined by the electrostatic boundary conditions. To evaluate Eq. (LABEL:eq:RashbaHamiltonianManyBody), we assume a perfectly flat TMDC monolayer with a $z$-dependent dielectric environment $\epsilon(z)$ without any vacancies Latini et al. (2015). Consequently, if we set $\epsilon_{1}=\epsilon_{2}$, i.e., equal substrate and superstrate material in Fig. 1, the field in Eq. (10) vanishes:
@@ III Excitonic Rashba Hamiltonian
Here, $\alpha_{\text{BR},v/c}^{\xi}$ is the Rashba coupling constant for holes/electrons at high-symmetry point $\xi$ Kormányos et al. (2014), $\varphi_{\nu,\mathbf{q}}^{\xi,\xi^{\prime},s,s^{\prime}}$ is the excitonic wave function as a solution of the Wannier equation, $\mathbf{q}$ is the relative momentum and $\alpha$, $\beta$ are the effective-mass ratios of the corresponding exciton configuration Katsch et al. (2018). We rewrote Eq. (15) and Eq. (16) in a way that they can address a more general situation compared to the dielectric environment only: $E_{z,\mathbf{q}}^{h/e,\text{tot}}$ can account for the total electric field having internal (local) contributions due to the dielectric environment and external contributions from, e.g., electrical gating, cf. Fig. 4. This way, the excitonic Rashba Hamiltonian in Eq. (LABEL:eq:RashbaHamiltonianIntrinsicExciton) combines two physically distinct fields, a local electric field $E_{z,\mathbf{q}}^{\text{loc}}$ due to many-particle interactions and an arbitrary in-plane spatially homogeneous external electric field $E_{z}^{\text{ext}}$, in one consistent description, which is described by a total electric field $E_{z,\mathbf{q}}^{h/e,\text{tot}}$ entering Eq. (15) and Eq. (16):
@@ V Spin Relaxation in Monolayer MoSe2
Figure 7: Dispersions of the spin-bright $E_{\mathbf{Q}}^{K,K,\uparrow,\uparrow}$ (red solid line) and spin-dark $E_{\mathbf{Q}}^{K,K,\uparrow,\downarrow}$ (blue solid line) excitons at the $K$ valley in a MoSe2 monolayer on a SiO2 substrate, cf. also Fig. 9. The green color gradient denotes the degree of the electron spin hybridization via a local electric field. Grey solid lines denote Boltzmann distributions of the spin-like excitonic occupation at various temperatures.
@@ 77 K 150 K 300 K
However, care should be taken in comparing specific numbers to distinct experiments, as a specific dielectric environment has been chosen for the calculation. Also, the bright-dark splitting plays a critical role. To illustrate this, we have also performed numerical simulations for a MoS2 monolayer for the same dielectric environment. The extracted spin relaxation times are displayed in Tab. 1 and are 5 ns at 77 K, which speed up to 25 ps at 300 K. Compared to a MoSe2 monolayer, they are orders of magnitude slower, which is a direct consequence of the larger bright-dark splitting of 14 meV Robert et al. (2020) compared to 1.45 meV in MoSe2 rendering the spin hybridization much less efficient. Other monolayer materials such as WSe2 or WSe2 behave similar, as their bright-dark splitting is even larger (40–55 meV Molas et al. (2019); Mapara et al. (2022); Molas et al. (2017); Zhou et al. (2017)). In these materials, other spin relaxation mechanisms such as spin-nonconserving scattering via chiral phonons Lagarde et al. (2024); Chan et al. (2025); Zhang and Cao (2022) are expected to dominate over a Rashba spin-orbit interaction induced by spatial asymmetries of the dielectric environment. However, in TMDC heterostructures such as MoSe2/MoS2 bilayers, intra-exciton fields due to charge-separated interlayer exciton densities can induce a significant many-body Rashba effect Mittenzwey et al. (2025). Thus, the exact choice of the material and dielectric environment conditions determine, which spin relaxation mechanism is dominant.
@@ VI Conclusion
In this work, we discovered a Rashba single spin-flip mechanism due to local out-of-plane electric fields induced by spatial asymmetries in the dielectric environment. We developed an excitonic many-body description, explicitly calculated the local electric fields for a given dielectric environment and quantified the Rashba coupling strength. For a monolayer MoSe2, we demonstrated very efficient spin relaxation on a sub-picosecond (150 K) to a hundred-femtosecond (300 K) timescale via a phonon-assisted single electron spin-flip channel.
@@ Acknowledgements.
H.M. and A.K. acknowledge financial support by the Deutsche Forschungsgemeinschaft (DFG) through Project KN 427/11-2 Project No. 420760124. A.K. acknowledges financial support by the DFG through Project KN 427/15-1, Project No. 556436549. We thank M. Snoeken for his help. H.M. thanks M. Selig for inspiration.

### v2

@@ Abstract
We propose a pair spin-orbit interaction (PSOI) mechanism by establishing a mesoscopic many-particle Rashba Hamiltonian. In lowest order, this Hamiltonian self-consistently describes exciton spin relaxation in monolayer transition metal dichalcogenides (TMDC) due to local electric fields caused by spatial asymmetries in the dielectric environment. For a monolayer MoSe2 on a SiO2 substrate above 77 K showing a meV bright-dark splitting, the local electric field causes fast intravalley spin relaxation on a sub-picosecond timescale, whereas it is negligible for other TMDCs with larger bright-dark splitting.
@@ I Introduction
Atomically thin semiconductors Wang et al. (2018a) host a broad variety of incoherent exciton scattering mechanisms. Based on the valley-selective optical selection rules, next to spin-conserving exciton-phonon scattering Selig et al. (2018); Brem et al. (2020); Wallauer et al. (2021); Katzer et al. (2023), Coulomb scattering mechanisms such as spin-conserving Dexter interaction Berghäuser et al. (2018) and spin-nonconserving intra- and intervalley exchange Guo et al. (2019); Maialle et al. (1993); Yu and Wu (2014); Selig et al. (2019, 2020); Kwong et al. (2021) via double spin flips are relevant. Moreover, single spin flips can occur via, e.g., Rashba interaction. Rashba spin-orbit interaction (SOI) can already affect linear optical spectra of a spin-forbidden mixture of $s$-$p$ orbital excitons in monolayers Cao et al. (2024, 2025). However, these signatures are very weak and have not been measured until now. For $s$-orbital excitonic states, which are usually measured in absorption or photoluminescence due to their large optical oscillator strength, no brightening of spin-dark excitonic states in out-of-plane electric fields can occur, since Rashba spin coupling vanishes at zero center-of-mass (COM) momentum within the light cone Mittenzwey et al. (2025). Thus, Rashba spin coupling is only relevant for incoherent $s$-orbital excitons occupying COM momenta well beyond the light cone. In contrast to Coulomb-exchange-induced double spin flips, single spin flip scattering mechanisms are less understood, as they are often included via external parameters Ochoa and Roldán (2013); Dery and Song (2015); Song and Dery (2013); Yang et al. (2020) or resulting from a complex band structure Wang et al. (2018b); Molina-Sánchez et al. (2017). Recently, chiral phonons have been identified as an important single spin-flip mechanism Lagarde et al. (2024); Chan et al. (2025); Zhang and Cao (2022). Besides the one-body SOI, it has been shown, that many-body-, Breit- or pair spin-orbit interaction (PSOI) due to self-induced Coulomb fields also occurs Breit (1929); Yi and Sheng (2000); Gindikin and Sablikov (2018, 2022); Liu and Principi (2024); Rajagopal and Mochena (1998), which completes the electron-electron interaction picture. In particular, such process can induce a superconducting phase in thin films under suitable conditions Gindikin and Kamenev (2025).
In this work, we derive a many-body Rashba PSOI mechanism Winkler (2003); Bychkov and Rashba (1984); Kormányos et al. (2014); Gindikin and Sablikov (2022); Chang (1993):
@@
which results from self-consistently described out-of-plane electric fields $\mathbf{E}(\mathbf{r},z)$, induced by the dielectric environment and examine its consequences for excitons in atomically thin semiconductors. Here, $\bm{\sigma}=\begin{pmatrix}\sigma_{x}&\sigma_{y}&\sigma_{z}\end{pmatrix}^{\top}$ is the vector of the Pauli matrices $\sigma_{i}$, $\mathbf{p}=\frac{\hbar}{\mathrm{i}}\nabla$ is the momentum operator and $\alpha_{\text{BR}}$ is the Rashba coupling constant, which, in solids, is greatly enhanced compared to the atomic constant in vacuum Kormányos et al. (2014); Winkler (2003).
Starting from Eq. (1), we derive a second-quantized description of the Rashba interaction induced by the optically induced (image) charge density due to spatial asymmetries in the dielectric environment of an atomically thin semiconductor (Sec. II), which is subsequently transformed into an excitonic description (Sec. III). In combination with exciton-phonon interaction generating non-zero COM-momentum excitons (Sec. IV), the described PSOI leads to a phonon-assisted exciton spin relaxation. We illustrate the new mechanism for a monolayer MoSe2 and quantify its impact on the incoherent exciton dynamics via microscopic calculations (Sec. V).
@@
$\begin{split}&\hat{H}_{\text{BR}}=\sum_{\begin{subarray}{c}\mathbf{k},\mathbf{k}^{\prime},\mathbf{q},\\ \lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4},\\ \xi,\xi^{\prime},\xi^{\prime\prime},s,s^{\prime}\end{subarray}}E_{z,\mathbf{q}+\mathbf{K}^{\xi^{\prime}}-\mathbf{K}^{\xi}}^{\text{loc}}F_{\lambda_{2},\lambda_{3},\mathbf{k}^{\prime}-\mathbf{q},\mathbf{k}^{\prime}}^{\xi^{\prime\prime}+\xi-\xi^{\prime},\xi^{\prime\prime},s^{\prime},s^{\prime}}\\ &\,\quad\times S_{\mathbf{k}+\mathbf{q},\mathbf{k}}^{\lambda_{1},\lambda_{4},\xi^{\prime},\xi,\bar{s},s}{\hat{a}^{\dagger}}{\vphantom{a}}_{\lambda_{1},\mathbf{k}+\mathbf{q}}^{\xi^{\prime},\bar{s}}{\hat{a}^{\dagger}}{\vphantom{a}}_{\lambda_{2},\mathbf{k}^{\prime}-\mathbf{q}}^{\xi^{\prime\prime}+\xi-\xi^{\prime},s^{\prime}}{\hat{a}^{\vphantom{\dagger}}}{\vphantom{a}}_{\lambda_{3},\mathbf{k}^{\prime}}^{\xi^{\prime\prime},s^{\prime}}{\hat{a}^{\vphantom{\dagger}}}{\vphantom{a}}_{\lambda_{4},\mathbf{k}}^{\xi,s}.\end{split}$ (9)
@@
where $\epsilon_{0}$ is the vacuum permittivity and $\epsilon_{s,\bot}$ is the relative out-of-plane dielectric constant of the thin semiconductor. The unitless functions $g_{\mathbf{q}}$ and $f_{\mathbf{q}}$, which encode the momentum dependence determined by the dielectric structure, are given in Eq. (41) and Eq. (42) in App. B, respectively. The electric field in Eq. (10) stems solely from the homogeneous solution of the Poisson equation, which arises due to image charges determined by the electrostatic boundary conditions.
To evaluate Eq. (LABEL:eq:RashbaHamiltonianManyBody), we assume a perfectly flat TMDC monolayer with a $z$-dependent dielectric environment $\epsilon(z)$ without any vacancies Latini et al. (2015). Consequently, if we set $\epsilon_{1}=\epsilon_{2}$, i.e., equal substrate and superstrate material in Fig. 1, the field in Eq. (10) vanishes:
@@ III Excitonic Rashba Hamiltonian
Here, $\alpha_{\text{BR},v/c}^{\xi}$ is the Rashba coupling constant for holes/electrons at high-symmetry point $\xi$ Kormányos et al. (2014), $\varphi_{\nu,\mathbf{q}}^{\xi,\xi^{\prime},s,s^{\prime}}$ is the excitonic wave function as a solution of the Wannier equation, $\mathbf{q}$ is the relative momentum and $\alpha$, $\beta$ are the effective-mass ratios of the corresponding exciton configuration Katsch et al. (2018). We rewrote Eq. (15) and Eq. (16) in a way that they can address a more general situation compared to the dielectric environment only: $E_{z,\mathbf{q}}^{h/e,\text{tot}}$ can account for the total electric field having internal (local) contributions due to the dielectric environment and external contributions from, e.g., electrical gating, cf. Fig. 4. This way, the excitonic Rashba Hamiltonian in Eq. (LABEL:eq:RashbaHamiltonianIntrinsicExciton) combines two physically distinct fields, a local electric field $E_{z,\mathbf{q}}^{\text{loc}}$ due to many-particle interactions inducing PSOI and an arbitrary in-plane spatially homogeneous external electric field $E_{z}^{\text{ext}}$ inducing SOI, in one consistent description, which is described by a total electric field $E_{z,\mathbf{q}}^{h/e,\text{tot}}$ entering Eq. (15) and Eq. (16):
@@ V Spin Relaxation in Monolayer MoSe2
Figure 7: Dispersions of the spin-bright $E_{\mathbf{Q}}^{K,K,\uparrow,\uparrow}$ (red solid line) and spin-dark $E_{\mathbf{Q}}^{K,K,\uparrow,\downarrow}$ (blue solid line) excitons at the $K$ valley in a MoSe2 monolayer on a SiO2 substrate, cf. also Fig. 9. The green color gradient denotes the degree of the electron spin hybridization via a local electric field. Grey solid lines denote Boltzmann distributions of the spin-bright excitonic occupation at various temperatures.
@@ 77 K 150 K 300 K
However, care should be taken in comparing specific numbers to distinct experiments, as a specific dielectric environment has been chosen for the calculation. Also, the bright-dark splitting plays a critical role. To illustrate this, we have also performed numerical simulations for a MoS2 monolayer for the same dielectric environment. The extracted spin relaxation times are displayed in Tab. 1 and are 5 ns at 77 K, which speed up to 25 ps at 300 K. Compared to a MoSe2 monolayer, they are orders of magnitude slower, which is a direct consequence of the larger bright-dark splitting of 14 meV Robert et al. (2020) compared to 1.45 meV in MoSe2 rendering the spin hybridization much less efficient. Other monolayer materials such as WSe2 or WSe2 behave similar, as their bright-dark splitting is even larger (40–55 meV Molas et al. (2019); Mapara et al. (2022); Molas et al. (2017); Zhou et al. (2017)). In these materials, other spin relaxation mechanisms such as spin-nonconserving scattering via chiral phonons Lagarde et al. (2024); Chan et al. (2025); Zhang and Cao (2022) are expected to dominate over a Rashba PSOI induced by spatial asymmetries of the dielectric environment. However, in TMDC heterostructures such as MoSe2/MoS2 bilayers, intra-exciton fields due to charge-separated interlayer exciton densities can induce a significant many-body Rashba effect Mittenzwey et al. (2025). Thus, the exact choice of the material and dielectric environment conditions determine, which spin relaxation mechanism is dominant.
@@ VI Conclusion
In this work, we established a Rashba single spin-flip mechanism due to local out-of-plane electric fields self-consistently induced by spatial asymmetries in the dielectric environment in atomically thin semiconductors. We developed an excitonic many-body description, explicitly calculated the local electric fields for a given dielectric environment and quantified the Rashba coupling strength. For a monolayer MoSe2, we demonstrated very efficient spin relaxation on a sub-picosecond (150 K) to a hundred-femtosecond (300 K) timescale via a phonon-assisted single electron spin-flip channel.
@@ Acknowledgements.
H.M. and A.K. acknowledge financial support by the Deutsche Forschungsgemeinschaft (DFG) through Project KN 427/11-2 Project No. 420760124. A.K. acknowledges financial support by the DFG through Project KN 427/15-1, Project No. 556436549. We thank Michiel Snoeken for his help. H.M. thanks Malte Selig for inspiration.
Data Availability Statement
The data supporting the findings of this work are available from the authors within reasonable request.